find the determinate of the matrix. Book shows below is 1 set or ()
(8+4a 4)
(5+3a 3)

Answer by venugopalramana(3286)

find the determinate of the matrix. Book shows below is 1 set or ()
(8+4a 4)
(5+3a 3)
D=(8+4A)*3-4(5+3A)=24+12A-20-12A=4